OpenStudy (anonymous):

i ams truggling with this graph can someone please help me? What is the possible discriminant of the graph?

OpenStudy (anonymous):

file attached below

OpenStudy (anonymous):

@austinL

OpenStudy (anonymous):

@sumi29

OpenStudy (anonymous):

@Peasant1 @julian25

OpenStudy (anonymous):

theses are my choices: –15 0 17 25 i think its -15 but i am not sure

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

It has two real roots, so the discriminant has to be something positive...

OpenStudy (anonymous):

if the discrimiant is negative then u will have complex solutions

OpenStudy (anonymous):

oh ok it wouldn't be zero so it is either 25 or 17 so by looking at the graph i am thinking it will be 17

OpenStudy (anonymous):

u r right

terenzreignz (terenzreignz):

I'm a bit uncertain in that regard, but I'd go for 17 too...

OpenStudy (anonymous):

it cant be 25 cuz 25 =5^2 and that meants integres soltuons like 1, 2, 3 ,4....

terenzreignz (terenzreignz):

No, it means rational solutions, but not integer solutions @julian25

OpenStudy (anonymous):

ok thanks for explaining this to me i struggle with graphs

OpenStudy (anonymous):

ohhh yes

terenzreignz (terenzreignz):

I mean, not *necessarily* integer solutions. That's why I'm still uncertain :D

OpenStudy (anonymous):

but u have irrational solutions there

terenzreignz (terenzreignz):

We don't know that, LOL the graph doesn't say anything :D

OpenStudy (whpalmer4):

This is a parabola, from the looks of it. Vertex will be halfway between the two crossings of the x-axis, at \(x =\frac{1}{2}(-4.56155-0.43845) = 2.5\) Equation of the parabola will be \[y = k(x + 4.56155)(x+0.43845)\]\[y = k(x^2+5x+2)\]with a suitable value of \(k\) to run it through whatever the y-coordinate of the vertex is. Doesn't matter for our purposes. Discriminant is \(b^2-4ac\) and we have \(a = 1, ~b= 5, ~c =2\)